Determine the dimensions of the right circular cylinder of

greatest volume that can be inscribed in a right circular cone of

radius 6 cm and height 9 cm.


1
Expert's answer
2022-05-11T17:28:30-0400

Using similar trianglr

Let h and r be the height and radius of cylinder respectively

Volume of cylinde:

"V=\pi r^2h"

"\frac{9}{6}=\frac{h}{6-r}"

"h=\frac{54-9r}{6}"

Put the value of the height into the formula for the volume of cylinder.

"V=\pi \frac{54-9r}{6}r^2=\frac{\pi*54*r^2}{6}-\frac{\pi*9r*r^2}{6}=9\pi*r^2-\frac{\pi*3r^3}{2}
V’=18\pi*r-4,5\pi*r^2"

"V’=18\pi*r-4,5\pi*r^2"

"V\u2019=0"

"18\pi*r-4,5\pi*r^2=0"

"9\pi*r(2-0,5r)=0"

"r=0"

and

"r=4"

"h=\frac{54-9*4}{6}=3"

"V=\pi*3*4^2=48\pi(cm^3)"